OpenStudy (anonymous):

Simplify ((2x^-2y^4)/(z^-5))^-3/((2xy^3)/(z^-2))

OpenStudy (anonymous):

\[\left[\begin{matrix}2x^{-2}y ^{4} \\ z ^{-5} \end{matrix}\right]^{-3} \frac{ 2xy ^{3} }{ z ^{-2} }\]